# The Multi-Range Theory of Translation Quality Measurement: MQM scoring models and Statistical Quality Control

### v3

@@ 7.3 The Non-Linear MQM Scoring Model
8 Discussion
@@
In this paper we covered a wide scope of sample sizes and different approaches and scoring models. It can be said that this paper represents the Unified Theory of Translation Quality Measurement which explains most use cases of translation quality measurement.
The FULL MQM Error Typology with Calibrated and non-linear scoring is a toolset which allows the reproduction of many different known proprietary metrics.
9 Future Work
@@
Acknowledgements
LH and GN are grateful for the grant “Integrating hospital outpatient letters into the healthcare data space” (EP/V047949/1; funder: UKRI/EPSRC).
Appendix A Appendix
A.1 Highlighted Scorecards and Model Parameters
@@
A.1.1 The MQM Metric Deployment Process (use case)

### v4

@@ 7.1 Non-Linear error tolerance – what it may look like?
8 Conclusion and Future Work
In this paper we covered a wide scope of sample sizes and different approaches and scoring models. It can be said that this paper represents the Unified Theory of Translation Quality Measurement which explains most use cases of translation quality measurement. The FULL MQM Error Typology with Calibrated and non-linear scoring is a toolset which allows the reproduction of many different known proprietary metrics.
Appendix A Parameters of the Evaluation Quality Metric
A.1 Calculation Values
The sample scorecard in Figure 8 shows the results of applying one possible metric to a specific evaluation task for a sample of segments (the size of which is stated in the Evaluation Word Count). Rows 29–35 list the selected Error Types – in this case the seven high-level error dimensions of MQM Core. The following values highlighted in Figure 8 play major roles in designing translation quality evaluation models. The abbreviations listed below are sometimes used when discussing formal equations.
A.1.1 Evaluation Word Count (EWC)
A.1.2 Reference Word Count (RWC)
A.2 Maximum Score Value (MSV)
A.2.1 Passing Threshold (PT)
A.2.2 Defined Passing Interval (DPI)
A.2.3 Final Quality Rating (Pass/Fail)
A.2.4 Error Type Weight (ETW)
A.3 Error Annotation Values
A.3.1 Error Type Number (ET No)
A.3.2 Error Severity Level
A.3.3 Severity Penalty Multiplier
A.4 Scoring Models Parameters
A.4.1 Error Count (EC)
A.4.2 Error Type Penalty Total (ETPT)
A.4.3 Absolute Penalty Total (APT)
A.4.4 Per-Word Penalty Total (PWPT)
A.4.5 Normed Penalty Total (NPT)
A.4.6 Quality Score (QS)
A.5 Calculating the Linear Quality Score
A.5.1 Quality Score without Calibration (Raw Score)
A.5.2 Quality Score with Calibration
The calibration process acts like a magnifying glass for viewing the otherwise very small or inconsistent acceptance ranges close to 100. This approach makes the quality rating easier to use and understand, highlighting differences in translation quality for evaluated texts more clearly.
A.6 Score Calculations
A.6.1 Calculating the raw quality scores
Calculating scores without calibration uses the steps shown in Figure 2. See the Appendix: Scoring Model Parameters for a list of all parameters and their abbreviations.
A.6.2 Calculating the quality scores with calibration
The scoring formula for calculating the quality score with calibration works with the standard calculation values, such as Evaluation Word Count, Absolute Penalty Total and Normed Penalty Total. However, a few additional values and parameters have to be defined. These are used to pre-define the specified acceptance criteria (the error tolerance) and to link these criteria to a scale that should be understandable or appropriate for all stakeholders. The following values used in the score calculation above are pivotal for a score calibrated with respect to a predefined Passing Threshold.
Scaling Factor (SF) Parameter to scale the Acceptable Penalty Points (APP) for the reference word count across the Defined Passing Interval (DPI). Let’s consider Figure LABEL:mqm2.0-scorecard-scaling-factor. On the top Raw Scale the raw passing threshold is 98, which means that a maximum of 20 raw penalty points are allowed on a sample of 1000 words. On a Calibrated Score scale (bottom) the Defined Passing Interval (DPI) is 15 (upward from 85 to 100). Therefore, the raw Passing Interval scales down from 20 on the raw scale to 15 on the Calibrated Scale. The Scaling Parameter for this new value will be 15/20=0.75. On the Raw Scale, the raw score is 98.44, which means that NPT=15.6 (the error density of a sample is equivalent to 15.6 errors on 1000 words). The trick to the Calibration is that what is scaled is not a passing interval, but rather, NPT. We multiply raw NPT of 15.6 by a scaling factor 15/20=0.75, which resolves in NPT = 11.7 on the calibrated scale. Therefore, the calibrated score will be 100-11.7=88.3 (bottom scale).
Appendix B Further Discussion
@@
Appendix C Highlighted Scorecards and Model Parameters
@@
Appendix D The MQM Metric Deployment Process (use case)